Circular No. 8062

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003X, 2003Y, AND 2003Z Three apparent new supernovae have been discovered on unfiltered

CCD images, as tabulated below. SN 2003X was reported by T. Boles (cf. IAUC 8059); SN 2003Y was found via LOTOSS/KAIT (reported by B. Beutler and W. Li; cf. IAUC 8061); and SN 2003Z was found by Y. L. Qiu and J. Y. Hu, Beijing Astronomical Observatory (BAO), with the BAO 0.6-m telescope (forwarded by W. Li).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003X Jan. 28.25 18^{h}12^{m}05.^{s}67 +29^{o}09′17′′.3 17.0 1′′.3 E, 4′′.2 S
2003Y Jan. 29.3  8 54 34.60 +57 10 19.8 17.4 3′′.7 W, 18′′.9 N
2003Z Jan. 29.7  9 07 32.46 +60 29 17.5 16.7 8′′.4 W, 31′′.0 N

Additional approximate magnitude estimates: SN 2003X in UGC 11151, 2002 Oct. 18 UT, [19.0 (Boles); 2003 Jan. 30.166, 17.0 (Boles); 30.529, 16.7 (Boles, from image by M. Schwartz, Cottage Grove, OR). SN 2003Y in IC 522, 2002 Dec. 25.3, [19.0 (KAIT); 2003 Jan. 30.3, 17.2 (KAIT). SN 2003Z in NGC 2742, Jan. 20.7, [18.1 (BAO); 30.4, 16.5 (KAIT).

COMET C/2002 X5 (KUDO-FUJIKAWA) R. W. Russell, D. K. Lynch, and D. L. Kim, The Aerospace Corpora-

tion; M. L. Sitko, University of Cincinnati; and W. Golisch, NASA Infrared Telescope Facility (IRTF), report that spectroscopy of comet C/2002 X5 was obtained on Jan. 9.1 and 10.05–10.14 UT at 1.5–3.4 airmasses with the Aerospace Broadband Array Spectrograph System at the IRTF 3-m telescope. They observed a relatively smooth thermal-emission spectrum, over most of the spectral range covered (3–14 μm), whose shape resembled a blackbody near or slightly above the blackbody radiative equilibrium temperature of 340 K. A silicate emission feature, if present, would be no more than ∼ 15 percent of the continuum. A 3.4-μm C-H feature may be present in emission. Narrowband magnitudes (each ± 0.05) on Jan. 9.1 in a 3′′.5diameter aperture were L [3.5 μm] = 8.3, M [4.5 μm] = 5.8, and N [10.2 μm] = 1.7. The emission at 10.2 μm was extended at least 18′′ east-west and 10′′ north-south, using a 29′′ north-south chop. The beam only needed to be moved ∼ 3′′ east-west to get to the half-power location, suggesting that the infrared emission was concentrated on the peak with an extended coma component. On Jan. 10, the spectrum was essentially unchanged, though perhaps 5 percent brighter.

2003 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green